# MF-LAL: Drug Compound Generation Using Multi-Fidelity Latent Space Active Learning

### v1

@@ 3.3 Generating molecules with high property scores
2:procedure generateHighScoringCompounds($k$, $M$, $\beta$)
@@
5:              initialize $\mathbf{z}^{(i)}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
@@
Our goal is to generate compounds at fidelity $k$ that maximize some generation objective. To accomplish this, we perform gradient-based optimization to find a point $\mathbf{z}_{k}$ in the $k$th latent space that maximizes the generation objective, and then decode $\mathbf{z}_{k}$ into a molecule using $p_{\theta_{k}}$. For our generation objective, we do not want to simply maximize $\hat{f}_{k}$, but instead the upper confidence bound (Eq. 2) to ensure exploration during active learning. In addition, we also introduce a novel likelihood-based term to the generation objective that encourages the model to only sample compounds at higher fidelities that also scored well at the lower fidelities. Specifically, when generating a molecule at fidelity $k$, we maximize the likelihood that the molecule would also be generated at fidelity $k-1$ with a high property score. This additional term greatly restricts the area of the chemical space explored by the high fidelity oracles, reducing the computational cost wasted on non-promising areas and making the use of high-cost oracles feasible. It also means the higher fidelity latent spaces encode a more limited distribution of compounds, improving the quality of samples generated from those latent spaces. Indeed, we show that the likelihood term is critical for strong performance (Table 2).
To compute the likelihood of a point $\mathbf{z}_{k}$ at fidelity $k$, we first generate a set of $M$ high-scoring compounds at fidelity $k-1$. Next, we map those points to a sum of Gaussians in the $k$th latent space using $h_{\xi_{k-1}}$, giving us a set of $M$ parameters $\{(\mu^{(j)}_{k},\sigma^{(j)}_{k})\}_{j=1}^{M}$. We then measure the likelihood of point $\mathbf{z}_{k}$ in the generated sum of Gaussians distribution. This guarantees that the compounds generated in latent space $k$ are also likely to have been generated in $k-1$ with high scores. Thus, we effectively reduce the size of the chemical space that must be explored at fidelity $k$ to only compounds that have already shown promise at the lower fidelities. Mathematically, maximizing the likelihood of a point $\mathbf{z}^{(i)}_{k}$ in latent space $k$ is equivalent to maximizing the probability density function evaluated at that point:
The full generation algorithm is detailed in Algorithm 2. In our implementation, we vectorize optimization across all $M$ latent space points simultaneously. In order to encourage diversity in generated compounds, we also add a term to the generation objective that measures the average pairwise cosine similarity between the $M$ points, and aim to minimize it:
where $s(A,B)$ is the cosine similarity between vectors $A$ and $B$.
@@ 4.1 Experimental setup
    Ensembled AutoDock4 ($f_{3}$) (Morris et al., 2009). Same as above, except we dock the compound into the binding pockets of eight BRD4/five c-MET cocrystal structures that were solved with different known ligands, and then take the minimum predicted energy. This ensemble approach is generally more accurate than using a single protein structure (Amaro et al., 2018).
@@
Each model is provided with an initial dataset of random ZINC250k (Irwin et al., 2012) compounds queried at each fidelity (see Appendix B for further details). To compare models, we run each in an active learning loop using a fixed computational budget of 7 days, and then generate 15 unique compounds at the highest fidelity predicted to have the best scores. We then run these compounds through ABFE and compare their scores. See Appendix C for additional results showing the oracle-predicted binding energy of the generated query compounds over the active learning process.
@@ 4.2 Baselines
  • SF-VAE (only ABFE / only docking) (Gómez-Bombarelli et al., 2018). Uses a simple single-fidelity GP as a surrogate model that is used to guide optimization in the latent space of a vanilla VAE, representing a simple single-fidelity approach. This consists of two separate baselines, one where the GP is trained only on ABFE data and one where it is trained on only docking data.
@@
  • VAE + 4x SF-GP. Uses a vanilla VAE model except with four independent GP surrogates, one for each fidelity, all using the single latent space as input. To be contrasted with MF-LAL, which uses multiple connected latent spaces instead of a single one.
  • VAE + MF-GP. Similar to above, except using a single multi-fidelity GP model (Wu et al., 2020) instead of four independent single-fidelity GP models.
@@
The first two baselines are single-fidelity methods, where we use both only ABFE and only docking as the single fidelity. The next four baselines, as well as MF-LAL, are multi-fidelity. Pocket2Mol does not utilize any oracle during generation. Evaluation of generated compounds from all baselines is done with ABFE.
@@ 4.3 Results
Table 1: Evaluation of generated compounds at highest fidelity. The mean and top 3 ABFE values are shown for 15 compounds sampled from each method after active learning for 7 days.
@@
Mean $\pm$ s.d. 1st 2nd 3rd Mean $\pm$ s.d. 1st 2nd 3rd
SF-VAE (only ABFE) -0.9 $\pm$ 2.7 -5.7 -2.9 -2.9 -1.2 $\pm$ 3.0 -4.4 -3.9 -3.1
SF-VAE (only docking) -3.1 $\pm$ 2.8 -6.1 -5.3 -4.8 -2.8 $\pm$ 3.4 -5.9 -5.8 -5.1
REINVENT (only ABFE) -3.9 $\pm$ 3.4 -8.7 -8.3 -8.2 -2.9 $\pm$ 3.7 -6.5 -5.8 -5.1
REINVENT (only docking) -3.1 $\pm$ 4.9 -11.0 -6.2 -5.7 -2.6 $\pm$ 5.0 -8.0 -6.8 -5.9
VAE + 4x SF-GP -2.3 $\pm$ 3.1 -8.0 -5.5 -5.3 -1.8 $\pm$ 2.5 -6.3 -5.9 -5.1
VAE + MF-GP -1.3 $\pm$ 3.3 -4.9 -3.1 -2.0 -3.3 $\pm$ 2.9 -9.7 -7.7 -4.2
MF-AL-GFN -2.5 $\pm$ 2.2 -6.5 -5.8 -5.1 -3.1 $\pm$ 1.8 -5.5 -4.5 -4.1
MF-AL-PPO -2.8 $\pm$ 2.5 -9.2 -6.5 -5.2 -4.0 $\pm$ 2.4 -7.0 -6.6 -5.8
Pocket2Mol -4.8 $\pm$ 4.6 -9.8 -9.8 -9.0 -2.2 $\pm$ 4.2 -4.5 -3.9 -3.2
MF-LAL (ours) -6.3 $\pm$ 4.0 -12.0 -11.3 -10.2 -6.5 $\pm$ 3.2 -12.9 -7.7 -6.5
Figure 2: Visualization of MF-LAL generated molecules. The top 3 molecules and associated ABFE scores are shown for both the BRD4(2) (top row) and c-MET (bottom row) targets.
Table 1 reports the average and top 3 ABFE scores of 15 compounds generated by MF-LAL, as well as those generated by the baseline methods, following active learning for 7 days. We ran each method separately for two targets, BRD4(2) and c-MET. We filtered generated compounds such that all had QED (Bickerton et al., 2012) $>0.4$, SAscore (Ertl & Schuffenhauer, 2009) $<4$, and no rings with $\geq 7$ atoms. We also only allowed compounds that fit these criteria to be queried during active learning. We filtered compounds that did not meet the criteria after generation, instead of performing multi-objective optimization, because most generated compounds from single-objective optimization already had a QED/SAscore in the range of typical drug compounds (Bickerton et al., 2012; Ertl & Schuffenhauer, 2009).
We find that the average ABFE scores of the compounds generated by MF-LAL, as well as those of the top three compounds, are significantly better (lower kcal/mol) than the corresponding scores of compounds generated by the baseline methods for both targets. The difference in average ABFE score (predicted binding free energy) between MF-LAL and the top baseline is -1.5 kcal/mol for BRD4(2) and -2.5 kcal/mol for c-MET, which are significant margins. Thus, our method outperforms both single and multi-fidelity techniques, as well as a 3D structure-based method (Pocket2Mol, Peng et al. (2022)) that does not use any binding affinity oracle. This suggests MF-LAL is most capable at generating compounds with real-world activity, since binding free energies computed with ABFE are the gold standard prediction method. Note, too, that the multi-fidelity techniques other than MF-LAL performed mostly similarly to the single-fidelity methods. This suggests that successfully taking advantage of multiple fidelities requires an architecture which, like that used in MF-LAL, is tailored to generating compounds for multiple fidelity oracles.
The molecular structures of the top 3 compounds generated by MF-LAL for both targets are shown in Figure 2. The compounds are structurally diverse, and none of them have close similars in the training set or in large datasets like PubChem (Kim et al., 2023). This shows the ability of MF-LAL to generate promising and novel structures. There is a small shared scaffold across 2 of the generated c-MET compounds consisting of a pyridazine, but this is not necessarily undesirable as it shows the model has found a high-property region of chemical space. The scaffold may be a promising starting point for development of empirical structure-activity relationships and lead optimization by medicinal chemists. Additionally, the pairwise Tanimoto similarity among the 15 compounds generated by MF-LAL is less than 0.2 for both targets, further indicating that our method generates a structurally diverse set.
Table 2: Ablations. The mean and top 3 ABFE-computed energies are shown for 15 compounds sampled from each method after active learning for 7 days.
@@
Mean $\pm$ s.d. 1st 2nd 3rd Mean $\pm$ s.d. 1st 2nd 3rd
MF-LAL -6.3 $\pm$ 4.0 -12.0 -11.3 -10.2 -6.5 $\pm$ 3.2 -12.9 -7.7 -6.5
-fid. 1 -6.1 $\pm$ 0.7 -7.7 -7.6 -7.4 -6.0 $\pm$ 1.1 -8.8 -7.0 -6.0
-fid. 2 -5.1 $\pm$ 2.0 -8.5 -6.5 -6.0 -5.2 $\pm$ 2.5 -8.0 -7.3 -6.1
-fid. 3 -4.2 $\pm$ 3.1 -9.2 -5.9 -5.7 -4.2 $\pm$ 3.5 -9.8 -7.1 -6.1
-fid. 4 -2.4 $\pm$ 3.2 -8.6 -4.3 -3.4 -3.1 $\pm$ 3.0 -7.6 -6.7 -5.1
w/o likelihood term -3.6 $\pm$ 4.4 -11.9 -10.7 -10.0 -3.8 $\pm$ 3.7 -10.9 -7.7 -6.3
Transformer enc/dec -6.2 $\pm$ 4.0 -11.5 -11.3 -9.8 -6.5 $\pm$ 2.9 -11.6 -7.6 -6.5
GCN enc/dec -5.9 $\pm$ 3.0 -10.9 -10.1 -9.3 -6.1 $\pm$ 4.6 -11.1 -7.5 -6.5
@@
Table 2 reports results from various ablations of the MF-LAL architecture. We experimented with removing each fidelity level individually, removing the likelihood term from the generation objective, and replacing the fully-connected encoder/decoder networks with a Transformer (Vaswani et al., 2017) and graph convolutional network (GCN, Kipf & Welling (2016)). The results show that all fidelities contribute to some degree to performance, although $f_{1}$ (the linear regression model) seems to add very little. Removing the likelihood term greatly reduced the performance of our method, showing that the approach of only querying compounds at higher fidelities that also scored well at lower fidelities is critical for strong performance. Replacing the fully-connected encoder/decoder with a Transformer had little effect on performance, so we used the simpler fully-connected version. Finally, changing the molecular representation to a graph and replacing the fully-connected encoder/decoder with GCNs (Kipf & Welling, 2016) resulted in slightly worse performance.

### v3

@@ 3.3 Generating molecules with high property scores
Our goal is to generate compounds at fidelity $k$ that maximize some generation objective. To accomplish this, we perform gradient-based optimization to find a point $\mathbf{z}^{(i)}_{k}$ in the $k$th latent space that maximizes the generation objective, and then decode $\mathbf{z}^{(i)}_{k}$ into a molecule using $p_{\theta_{k}}$. For our generation objective, we do not want to simply maximize $\hat{f}_{k}$, but instead the upper confidence bound (Eq. 1) to ensure exploration during active learning. In addition, we also introduce a novel likelihood-based term to the generation objective that encourages the model to only sample compounds at higher fidelities that also scored well at the lower fidelities. Specifically, when generating a molecule at fidelity $k$, we maximize the likelihood that the molecule would also be generated at fidelity $k-1$ with a high property score. This additional term greatly restricts the area of the chemical space explored by the high fidelity oracles, reducing the computational cost wasted on non-promising areas and making the use of high-cost oracles feasible. It also means the higher fidelity latent spaces encode a more limited distribution of compounds, improving the quality of samples generated from those latent spaces. Indeed, we show that the likelihood term is critical for strong performance (Table 3).
To compute the likelihood of a point $\mathbf{z}^{(i)}_{k}$ at fidelity $k$, we first generate a set of $M$ high-scoring compounds at fidelity $k-1$. Next, we map those points to a sum of Gaussians in the $k$th latent space using $h_{\xi_{k-1}}$, giving us a set of $M$ parameters $\{(\mu^{(j)}_{k},\sigma^{(j)}_{k})\}_{j=1}^{M}$. We then measure the likelihood of point $\mathbf{z}_{k}$ in the generated sum of Gaussians distribution (see Appendix A.1 for mathematical details). This guarantees that the compounds generated in latent space $k$ are also likely to have been generated in $k-1$ with high scores. Thus, we effectively reduce the size of the chemical space that must be explored at fidelity $k$ to only compounds that have already shown promise at the lower fidelities.
@@
2:procedure generateHighScoringMols($k$, $M$, $\beta$)
@@
5:              initialize $\mathbf{z}^{(i)}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ as a starting point for gradient-based optimization
@@
The full generation algorithm is detailed in Algorithm 2. In our implementation, we vectorize optimization across all $M$ latent space points simultaneously. In order to encourage diversity in generated compounds, we also add a term to the generation objective that measures the average pairwise cosine similarity between the $M$ points (Appendix A.1)
@@ 4.1 Experimental setup
    Ensembled AutoDock4 ($f_{3}$) (Morris et al., 2009). Same as above, except we dock the compound into the binding pockets of 8 BRD4/5 c-MET cocrystal structures that were solved with different known ligands, and then take the minimum predicted energy. This ensemble approach is generally more accurate than using a single protein structure (Amaro et al., 2018).
@@
Each model is provided with an initial dataset of random ZINC250k (Irwin et al., 2012) compounds queried at each fidelity (see Appendix B for further details). To compare models, we run each in an active learning loop using a fixed computational budget of 7 days, and then generate 15 unique compounds at the highest fidelity predicted to have the best scores. We then run these compounds through ABFE and compare their scores. Based on the real-world use case of our method, and following other works (Lee et al., 2023; Luo et al., 2021; Fu et al., 2022), we focus particularly on the properties of the top 3 generated compounds. This is because drug campaigns would take only the top compounds generated and use them as starting points for further optimization, so the binding affinities of the top compounds are the most relevant for measuring performance. See Appendix C for additional results showing the oracle-predicted binding energy of the generated query compounds over the active learning process, and the reconstruction accuracy for each fidelity decoder during training.
@@ 4.2 Baselines
  • SF-VAE (only ABFE / only docking) (Gómez-Bombarelli et al., 2018). Uses a simple single-fidelity GP as an activity surrogate model that is used to guide optimization in the latent space of a vanilla VAE, representing a simple single-fidelity approach. This consists of two separate baselines, one where the GP is trained only on ABFE data and one where it is trained on only docking data.
@@
  • RGA (only docking) (Fu et al., 2022). A genetic algorithm-based method that uses a 3D neural network on both the protein and ligand to prioritize mutations and crossovers. The docking score is used as the reward function, making it a single-fidelity method.
  • VAE + 4x SF-GP. Uses a vanilla VAE model for molecule generation and four independent GP surrogates for activity prediction, one for each fidelity, all using the single latent space as input. To be contrasted with MF-LAL, which uses multiple connected latent spaces instead of a single one.
  • VAE + MF-GP. Similar to above, except using a single multi-fidelity GP model activity surrogate (Wu et al., 2020) instead of four independent single-fidelity GP models.
@@
  • MF-GP + ZINC250k. Active learning using a multi-fidelity GP model (Wu et al., 2020) with a fixed candidate set consisting of ZINC250k compounds (Irwin et al., 2012). This baseline represents a non-query synthesis approach to multi-fidelity active learning.
@@
  • DecompDiff (Guan et al., 2023b). Same type of model as above, but makes additional improvements to the generation process by decomposing generation into scaffold and motif stages.
  • TAGMol (only docking) (Dorna et al., 2024). 3D structure-based diffusion model that conditions generation on the protein pocket. A property predictor is introduced to guide the diffusion process towards ligands that have high predicted affinity. In our case, the property predictor is a single-fidelity predictor of the docking score.
The first three baselines are single-fidelity methods, where we use both only ABFE and only docking as the single fidelity. The next five baselines, as well as MF-LAL, are multi-fidelity. Pocket2Mol and DecompDiff do not utilize any oracle during generation. Evaluation of generated compounds from all baselines is done with ABFE.
@@ 4.3 Results
Figure 3: Visualization of MF-LAL generated molecules. The top 3 molecules and associated ABFE scores are shown for both the BRD4(2) (top row) and c-MET (bottom row) targets. The compounds are diverse and relatively drug-like.
Table 1: Evaluation of 15 generated compounds at highest fidelity. The mean and top 3 ABFE values are shown for 15 compounds sampled from each method after active learning for 7 days. “# active scafs.” is the number of compounds (“scaffolds”) that are active ($<-8.2$ kcal/mol for BRD4(2) and $<-6.8$ for c-MET) and have $<0.4$ Tanimoto similarity to any other active compound.
@@
# active # active
Mean $\pm$ sd scafs. Count 1st 2nd 3rd Mean $\pm$ sd scafs. Count 1st 2nd 3rd
SF-VAE (only ABFE) -0.9 $\pm$ 2.7 0 15 -5.7 -2.9 -2.9 -1.2 $\pm$ 3.0 0 15 -4.4 -3.9 -3.1
SF-VAE (only docking) -3.1 $\pm$ 2.8 0 15 -6.1 -5.3 -4.8 -2.8 $\pm$ 3.4 0 15 -5.9 -5.8 -5.1
REINVENT (only ABFE) -3.9 $\pm$ 3.4 2 15 -8.7 -8.3 -8.2 -2.9 $\pm$ 3.7 0 15 -6.5 -5.8 -5.1
REINVENT (only docking) -3.1 $\pm$ 4.9 1 15 -11.0 -6.2 -5.7 -2.6 $\pm$ 5.0 1 15 -8.0 -6.8 -5.9
RGA (only docking) -3.1 $\pm$ 3.9 0 15 -7.8 -7.0 -6.8 -2.1 $\pm$ 3.0 0 15 -6.0 -5.5 -5.4
VAE + 4x SF-GP -2.3 $\pm$ 3.1 0 15 -8.0 -5.5 -5.3 -1.8 $\pm$ 2.5 0 15 -6.3 -5.9 -5.1
VAE + MF-GP -1.3 $\pm$ 3.3 0 15 -4.9 -3.1 -2.0 -3.3 $\pm$ 2.9 1 15 -9.7 -6.0 -4.2
MF-AL-GFN -2.5 $\pm$ 2.2 0 15 -6.5 -5.8 -5.1 -3.1 $\pm$ 1.8 0 15 -5.5 -4.5 -4.1
MF-AL-PPO -2.8 $\pm$ 2.5 1 15 -9.2 -6.5 -5.2 -4.2 $\pm$ 2.8 0 15 -6.6 -5.8 -5.5
MF-GP + ZINC250k -3.0 $\pm$ 2.9 0 15 -5.5 -4.7 -4.5 -2.9 $\pm$ 3.1 0 15 -6.3 -5.8 -5.0
Pocket2Mol -4.3 $\pm$ 3.8 1 15 -9.8 -8.7 -8.0 -2.2 $\pm$ 4.2 0 15 -4.5 -3.9 -3.2
DecompDiff -2.7 $\pm$ 4.0 1 15 -8.9 -8.1 -7.5 -1.9 $\pm$ 6.4 1 15 -8.0 -5.1 -2.7
TAGMol (only docking) -1.9 $\pm$ 3.0 0 15 -8.1 -7.6 -6.9 -3.5 $\pm$ 3.3 1 15 -7.0 -5.6 -5.1
MF-LAL (ours) -6.2 $\pm$ 3.9 6 15 -12.0 -10.2 -9.8 -6.7 $\pm$ 3.1 4 15 -12.9 -7.9 -7.7
Table 2: Evaluation of 40 generated compounds at highest fidelity. For MF-LAL and the most competitve baseline (Table 1), we ran ABFE on an additional 25 generated compounds for increased robustness. * refers to $p<0.05$.
@@
# active # active
Mean $\pm$ sd scafs. Count 1st 2nd 3rd Mean $\pm$ sd scafs. Count 1st 2nd 3rd
MF-AL-PPO -4.3 $\pm$ 2.6 1 40 -8.5 -6.6 -5.8
Pocket2Mol -4.6 $\pm$ 3.8 2 40 -9.8 -9.8 -9.3
MF-LAL (ours) -6.3* $\pm$ 3.7 8* 40 -12.0 -11.3 -10.2 -7.1* $\pm$ 3.0 6* 40 -13.9 -12.9 -7.9
Table 1 reports the average and top 3 ABFE scores of 15 compounds generated by MF-LAL, as well as those generated by the baseline methods, following active learning for 7 days. We ran each method separately for two targets, BRD4(2) and c-MET. For MF-LAL and the most competitive baseline for each target (evaluated by lowest mean ABFE score), we evaluated an additional 25 compounds (total 40) for improved robustness, which is shown in Table 2. We also report the number of active scaffolds, which means the number of generated compounds that are active and have less than 0.4 Tanimoto similarity to any other active compound. We defined active as $<-8.2$ kcal/mol ($<1\mu M$ activity) for BRD4(2) and $<-6.8$ kcal/mol ($<10\mu M$ activity) for c-MET based on the affinities of compounds from previous works that investigate these targets (Liu et al., 2007, 2017; Naguib et al., 2024). Specifically, Liu et al. (2017) consider BRD4(2) inhibitors active when they have submicromolar activity, and Naguib et al. (2024) state that “potent activity” against c-MET is achieved with a 12 $\mu$M inhibitor. The number of active scaffolds metric closely reflects the goal of early-stage drug discovery, which is to identify a diverse set of highly active compounds.
We filtered generated compounds such that all had QED (Bickerton et al., 2012) $>0.4$, SAscore (Ertl & Schuffenhauer, 2009) $<4$, and no rings with $\geq 7$ atoms. 55% of all compounds generated for BRD4(2), and 68% for c-MET, fulfilled these criteria. We also only allowed compounds that fit these criteria to be queried during active learning. We filtered compounds that did not meet the criteria after generation, instead of performing multi-objective optimization, because most generated compounds from single-objective optimization already had a QED/SAscore in the range of typical drug compounds (Bickerton et al., 2012; Ertl & Schuffenhauer, 2009). Following this filtering step, the mean QED of generated compounds for BRD4(2) (c-MET) was 0.59 (0.63), SAscore was 3.6 (3.5), and diversity (1 - mean pairwise Tanimoto similarity) was 0.81 (0.83).
Table 3: Ablations on model architecture. The mean and top 3 ABFE-computed energies are shown for 15 compounds sampled from each method after active learning for 7 days.
Method BRD4(2) ABFE (kcal/mol) c-MET ABFE (kcal/mol)
# active # active
Mean $\pm$ sd scafs. 1st 2nd 3rd Mean $\pm$ sd scafs. 1st 2nd 3rd
MF-LAL (ours) -6.2 $\pm$ 3.9 6 -12.0 -10.2 -9.8 -6.7 $\pm$ 3.1 4 -12.9 -7.9 -7.7
-fid. 1 -6.1 $\pm$ 0.7 0 -7.7 -7.6 -7.4 -6.0 $\pm$ 1.1 1 -8.8 -7.0 -6.0
-fid. 2 -5.1 $\pm$ 2.0 1 -8.5 -6.5 -6.0 -5.2 $\pm$ 2.5 2 -8.0 -7.3 -6.1
-fid. 3 -4.2 $\pm$ 3.1 1 -9.2 -5.9 -5.7 -4.2 $\pm$ 3.5 1 -9.8 -7.1 -6.1
-fid. 4 -2.4 $\pm$ 3.2 1 -8.6 -4.3 -3.4 -3.1 $\pm$ 3.0 1 -7.6 -6.7 -5.1
w/o likelihood term -3.4 $\pm$ 4.1 2 -11.9 -9.7 -9.0 -3.8 $\pm$ 3.7 2 -10.9 -7.7 -6.3
Transformer enc/dec -6.1 $\pm$ 3.8 3 -11.5 -9.9 -9.0 -6.5 $\pm$ 2.9 2 -11.6 -7.6 -6.5
GCN enc/dec -5.9 $\pm$ 3.0 2 -10.9 -10.1 -9.0 -6.1 $\pm$ 4.6 1 -11.1 -7.5 -6.5
We find that the average ABFE scores of the compounds generated by MF-LAL, as well as those of the top three compounds, are significantly better (lower kcal/mol) than the corresponding scores of compounds generated by the baseline methods for both targets. The difference in average ABFE score (predicted binding free energy) between MF-LAL and the top baseline is -1.6 kcal/mol for BRD4(2) ($p=0.03$) and -2.8 kcal/mol for c-MET ($p<0.001$), which are significant margins (see Appendix B.3 for details on the statistical tests). Additionally, MF-LAL generates significantly more active scaffolds than the most competitive baseline ($p<0.05$ for both targets), and the mean ABFE score of the top 3 compounds from MF-LAL is significantly lower than the mean ABFE score of the top 3 compounds from baseline methods ($p<0.05$ for both targets). Thus, our method outperforms both single and multi-fidelity techniques, as well as 3D structure-based methods (Pocket2Mol (Peng et al., 2022), DecompDiff (Guan et al., 2023b), and TAGMol (Dorna et al., 2024)) that do not use any binding affinity oracle. This suggests MF-LAL is most capable at generating compounds with real-world activity, since ABFE scoring is the gold standard prediction method. Note, too, that the multi-fidelity techniques other than MF-LAL performed mostly similarly to the single-fidelity methods. This suggests that successfully taking advantage of multiple fidelities requires an architecture which, like that used in MF-LAL, is tailored to generating compounds for multiple fidelity oracles.
The molecular structures of the top 3 compounds generated by MF-LAL for both targets are shown in Figure 3. The compounds are structurally diverse, and none of them have close similars in the training set or in large datasets like PubChem (Kim et al., 2023). This shows the ability of MF-LAL to generate promising and novel structures. There are some similar ring structures across the compounds generated for c-MET, but this is not necessarily undesirable as it shows the model has found a high-property region of chemical space. The scaffold may be a promising starting point for development of empirical structure-activity relationships and lead optimization by medicinal chemists. Additionally, the mean pairwise Tanimoto similarity among the 40 generated compounds is $<$ 0.2 for both targets, further indicating that our method generates a structurally diverse set.
@@
Table 3 reports results from various ablations of the MF-LAL architecture (using 15 compounds sampled from all methods). We experimented with removing each fidelity level individually, removing the likelihood term from the generation objective, and replacing the fully-connected encoder/decoder networks with a Transformer (Vaswani et al., 2017) and graph convolutional network (GCN for encoder and inner product decoder as described by Kipf & Welling (2016)). The results show that all fidelities contribute to performance. Removing the likelihood term greatly reduced the performance of our method, showing that the approach of only querying compounds at higher fidelities that also scored well at lower fidelities is critical for strong performance. Replacing the fully-connected encoder/decoder with a Transformer had little effect on performance, so we used the simpler fully-connected version. Finally, changing the molecular representation to a graph and replacing the fully-connected encoder/decoder with GCNs (Kipf & Welling, 2016) resulted in slightly worse performance.